An example of a problem is:
Find the value of y when x = -2
y= 2x2 – 3x +15


Answer by CharlesG2(834)

An example of a problem is:
Find the value of y when x = -2
y= 2x2 – 3x +15
y = 2 * (-2)^2 - 3(-2) + 15
y = 2 * 4 + 6 + 15
y = 8 + 6 + 15
y = 14 + 15
y = 29